# Adaptive Visual Scene Understanding: Incremental Scene Graph Generation

### v1

@@ A.2.3 Definitions of BWT and FWT Evaluation Metrics
To assess the influence that learning a task $t$ has on the performance of any previous tasks in CSEGG models, we also report Backward Transfer (BWT@K)(Lopez-Paz & Ranzato, 2017). BWT@K is defined as $BWT@K=\frac{1}{T-1}\sum_{i=1}^{T-1}R@K_{T,i}\>-\>R@K_{i,i}$, where $T$ denotes the total number of tasks in a learning scenario and $R@K_{i,j}$ denotes the continual learning model trained after task $i$ and tested in task $j$.
To assess the influence of previous tasks on the current task $t$ in CSEGG models, we also report Forward Transfer (FWT@K) (Lin et al., 2022). FWT is defined as $FWT@K=\frac{1}{T-1}\sum_{i=2}^{T}R@K_{i,i}\>-\>\overline{b@K}_{i,i}$, where $\overline{b@K}_{i,i}$ is the R@K for an independent model with random initialization trained in task $i$ and tested in task $i$.
A.3 Data Statistics
In this section, we provide various types of data statistics for all three learning scenarios. Specifically, we present statistics regarding the number of images, objects, and relationships involved in each task of each learning scenario. Additionally, we include data statistics pertaining to replay buffers of sizes 10%, 20%, and 100%.
A.3.1 Scenario 1 (S1): Relationship Incremental Learning
In this scenario, new relationship predicates $p_{k}$ are incrementally added in each task. The Learning Scenario 1 (S1) comprises five tasks, with each task consisting of 10 mutually exclusive relationships. Across all tasks, there is a common set of 150 objects present. Fig. S1 shows comprehensive data statistics for stages 1 and 2 over all the tasks.
As mentioned in Sec. 3.2, we skip the Stage 1 training in S1 and directly load pre-trained weights of DETR (Zhu et al. (2020)) for object detection on the entire training set of Visual Genome. Fig. S1(a) shows the distribution of objects in the entire training set of Visual Genome.
Fig. S1(b) displays the distribution of relationships during Training Stage 2 for each task in the Learning Scenario 1 (S1). Each task consists of 10 relationships that are mutually exclusive. Notably, a long tail pattern is observed in the distribution of relationships for each task. The legend of Fig. S1(b) indicates that there is a relatively uniform number of images present for each task in Stage 2 of the Learning Scenario 1 (S1).
The test sets for each task in S1 exhibit the distributions of the number of images and relationships that closely align with the distributions of the training sets depicted in Fig. S1(b).
A.3.2 Scenario 2 (S2): Scene Incremental Learning
In this learning scenario, new objects $O_{i}$ and new relationship predicates $p_{k}$ are incrementally introduced over tasks. The Learning Scenario 2 (S2) comprises of 2 tasks with the first task containing 100 objects and 40 relationships and the second task containing 25 objects and 5 relationships. Fig. S2 shows comprehensive data statistics pertaining to Learning Scenario 2 (S2).
As mentioned in Sec. 3.2, both the Stage 1 and Stage 2 training are present in S2. Fig. S2(a) presents the distribution of objects during Stage 1 for each task in the S2. The first task consists of 100 objects and the second task consists of 25 objects which are mutually exclusive. Notably, a long tail pattern is observed in the distribution of objects for each task. The legend of Fig. S2(a) indicates that there is a uniform number of images present for each task in Stage 1 of the S2.
Fig. S2(b) displays the distribution of relationships during Stage 2 for each task in the S2. The first task consists of 40 relationships and the second task consists of 5 relationships that are mutually exclusive. Notably, a long tail pattern is observed in the distribution of relationships for each task. The legend of Fig. S2(b) highlights that in Stage 2 of the S2, Task 1 has a noticeably higher number of images compared to Task 2. However, the number of relationship notations per class from Task 2 is higher than Task 1, since relationships from head classes are assigned to Task 2 (see Sec. 3.1).
The test sets for each task in S2 exhibit similar distributions in the number of images, objects, and relationships to the distributions of the training sets depicted in Fig. S2(a)(b).
A.3.3 Scenario 3 (S3): Scene Graph Generalization In Object Incremental Learning
In this learning scenario, there are a total of four tasks, with each task encompassing 30 distinct objects and 35 common relationships. Fig. S3 shows comprehensive data statistics pertaining to Learning Scenario 3 (S3).
As mentioned in Sec. 3.2, both Stage 1 and Stage 2 training are present in S3. Fig. S3(a) presents the distribution of objects during Stage 1 for each task in the S3. Each task consists of 30 objects which are mutually exclusive. Notably, a long tail pattern is observed in the distribution of objects for each task. The legend of Fig. S3(a) indicates that there is a relatively uniform number of images present for each task in Stage 1 of the S3. Fig. S3(b) displays the distribution of relationships during Stage 2 for each task in the S3. Each task consists of 35 relationships which are common over all the tasks. Notably, a long tail pattern is observed in the distribution of relationships for each task.
As mentioned in Sec. 3.1 of Scenario 3 (S3), a standalone generalization test set is curated for S3. This test set consists of 1942 images and features a completely different set of objects compared to those present in the training data. However, the test set maintains the same set of relationships as observed in the training data.
A.3.4 Replay Buffers
In our benchmarking process, we evaluate multiple replay baselines using a memory buffer with a fixed capacity to store a certain percentage (M) of images from the entire training dataset, along with their corresponding ground truth object and relationship notations specific to each learning scenario. We vary the value of M to be 10%, 20%, and 100%. As there could be multiple ground truth objects and relationships per image, we provide statistics on the number of ground truth notations stored in the memory buffer for each task in all the learning scenarios in this section. Specifically, we report the histograms for ground truth notations in the replay buffer for scenario 1 (Fig. S1), scenario 2 (Fig. S2), and scenario 3 (Fig. S3).
In general, over all three learning scenarios, we observe a long-tail distribution of notations per object or relationship class within each task. At stage 1, as we fix the memory capacity over tasks, the number of images stored in the memory buffer remains constant over tasks in most cases, except for memory capacity $M=100\%$. This is to maximize the diversity of information for replays. Note that although the number of images remain constant in the memory buffer, the number of images allocated to each task decreases given a fixed memory buffer capacity; and hence, we saw a decrease in the number of notations from previous tasks.
At stage 2, the images stored in stage 1 are carried over. However, as some images might not contain task-relevant relationship classes, the number of images used for training at stage 2 might vary over tasks. The notations stored in the memory buffer also follow the long-tail distributions.
In learning scenario 1, we reported the histograms of ground truth notations after two sampling techniques, LVIS and BLS, are applied in the memory buffers (Fig. S4). Both LVIS and BLS resampling methods aim to address the long-tail distribution issues by either reducing the prominence of head classes or over-sampling the object instances from the tail classes. Indeed, by comparing the histograms in Fig. S4(a)(b), we noticed such changes in histograms for tail and head classes. A similar trend in the class distribution per task emerges when applying long-tail distribution techniques to the exemplar dataset, as indicated in Fig.S4(c)(d).
Figure S1: Data Statistics for Learning Scenario 1 (S1): (a) Distribution of objects in the entire training set of Visual Genome during Stage 1. (b) Distribution of relationships during Stage 2 for each task in S1. (c) and (d) show distributions of relationships in the memory buffer for Replay 10% and Replay 100 % during Stage 2 for each task in S1. See the legend for the color codes of each task. The numbers in brackets in the legend in (b-d) denote the number of training images in the particular task. Zoom in to the figure to get the exact labels and the frequency associated with them.
Figure S2: Data Statistics for Learning Scenario 2 (S2): (a) Distribution of objects during Stage 1 for each task in S2. (b) Distribution of relationships during Stage 2 for each task in S2.(c) and (d) show distributions of objects in the memory buffer for Replay 10% and Replay 20% during Stage 1 for each task in S2. (e) and (f) show distributions of relationships in the memory buffer for Replay 10% and Replay 20% during Stage 2 for each task in S2. See the legend for the color codes of each task. The numbers in brackets in the legend in (a-f) denote the number of training images in the particular task. Zoom in to the figure to get the exact labels and the frequency associated with them.
Figure S3: Data Statistics for Learning Scenario 3 (S3): (a) Distribution of objects during Stage 1 for each task in S3. (b) Distribution of relationships during Stage 2 for each task in S3. (c) Distribution of objects in the memory buffer for Replay 10% during Stage 1 for each task in S3. (d) Distribution of relationships in the memory buffer for Replay 10% during Stage 2 for each task in S3. See the legend for the color codes of each task. The numbers in brackets in the legend in (a-d) denote the number of training images in the particular task. Zoom in to the figure to get the exact labels and the frequency associated with them.
Figure S4: Data Statistics for Learning Scenario 1 (S1) when sampling techniques on long-tailed distribution are applied: (a) Distribution of relationships during Stage 2 for each task in S1 when LVIS is applied. (b) Distribution of relationships in the memory buffer for Replay 10% during Stage 2 for each task in S1 when LVIS is applied.(c) Distribution of relationships during Stage 2 for each task in S1 when BLS is applied. (d) Distribution of relationships in the memory buffer for Replay 10% during Stage 2 for each task in S1 when BLS is applied. The numbers in brackets in the legend in (a-d) denote the number of training images in the particular task. Zoom in to the figure to get the exact labels and the frequency associated with them.
A.4 More Quantitative Result Analysis on Continual SGTR Methods
A.4.1 Results for K = 50, 100
Fig. S5 and S6 present the results for Avg.R@K and F@K with K values of 50 and 100, focusing on the Learning Scenarios 1 and 2. These results align with the findings discussed in Sec. 4.1, which considered Avg.R@20 and F@20. Notably, when K is increased to 50 and 100, both Avg.R and F exhibit higher absolute values compared to K=20 but the trends remain consistent. Fig. S7 presents the results of Gen.R@K for K=50,100 for Learning Scenario 3. Similar findings in Sec. 4.3 can be made here as well.
Figure S5: Results in AvgR@K and F@K over tasks in Learning Scenario 1 (S1) when K=50,100. See Fig. 5 for design conventions.
Figure S6: Results in AvgR@K and F@K over tasks in Learning Scenario 2 (S2) when K=50,100. See Fig. 5 for design conventions.
Figure S7: Results in Gen R@K over tasks in Learning Scenario 3 (S3) when K=50,100. See Fig. 6 for design conventions.
A.4.2 Mean Recall Results
Mean Recall@K (mR@K) calculates Recall@K for each relationship category independently and then reports their mean. This is a more fair evaluation metric when dealing with long tail distributions as it averages out the recall over all the relationship classes. In this section, we define mF@K and Avg.mR@K based on mR as described in Sec. 3.4. We report the results of mF@K and Avg.mR@K for learning scenarios 1 and 2 in Fig. S8 and Fig. S9. Note that the evaluations with mF@K and Avg.mR@K are not applicable in learning scenario 3. Though the absolute values of mF@K and Avg.mR@K are generally smaller than the values of F@K and Avg.R@K in Fig. 5, the same observations made in Sec. 4.1 can be applied here as well.
Figure S8: Results in Avg. mR@K and mF@K over tasks in Learning Scenario 1 (S1) when K=20,50,100. See Fig. 5 for design conventions.
Figure S9: Results in Avg. mR@K and mF@K over tasks in Learning Scenario 2 (S2) when K=20,50,100. See Fig. 5 for design conventions.
Figure S10: Results in Avg. mR@K and mF@K over tasks in Learning Scenario 1 (S1) when sampling techniques on long-tailed distribution are applied.. See Sec. 3.2 for the introduction to techniques used for long-tailed distributions. See Tab. 1 for design conventions.
A.4.3 Result Analysis on BWT@K and FWT@K
We provide the results of BWT@K and FWT@K in Fig. S11 for Learning Scenario 1. Consistent with the results of Avg.R@20 and F@20, we observe that among all methods, the naive method has the lowest FWT@20 and BWT@20 implying catastrophic forgetting. Once again, EWC outperforms the naive baseline by having higher FWT@20 and BWT@20 though the difference is small as it fails in longer task sequences. Although all the baselines have negative BWT@20 indicating forgetting, Replay(10%) yields a higher BWT@20 than PackNet, EWC and naive baseline pointing to a reduced level of forgetting. Replay(100%) exhibits the greatest BWT@20, suggesting the significance of revisiting previous samples to prevent forgetting and facilitate backward knowledge transfer. Similar to BWT@20, most of the baselines have negative FWT@20 implying that knowledge acquired from prior tasks slightly interferes with the learning process of the current task. In contrast, Replay(10%) shows positive FWT@20 indicating that knowledge acquired from prior tasks enhances the learning process of the current task. Surprisingly, while Replay(100%) excels in Avg.R@20, F@20, and BWT@20, Replay(10%) achieves a superior FWT@20. This suggests that Replay(100%) might not grasp the current task as effectively as Replay(10%), as it preserves the knowledge from previous tasks by replaying more old samples and forgetting less at the sacrifice of slow learning about the current task. Similar trends are observed and the same analysis can be applied in Learning Scenario 2 (Fig. S12).
Figure S11: Results of FWT (a) and BWT (b) over baselines in Learning Scenario 1 based on the SGTR backbone. See Sec. 3.3 for introduction to continual learning baselines. See Sec. 3.4 for explanations about evaluation metrics. The x-axis indicates the various baselines. The higher FWT@20 and BWT@20, the better.
Figure S12: Results of FWT (a) and BWT (b) over baselines in Learning Scenario 2 based on the SGTR backbone. See Sec. 3.3 for introduction to continual learning baselines. See Sec. 3.4 for explanations about evaluation metrics. The x-axis indicates the various baselines. The higher FWT@20 and BWT@20, the better.
A.4.4 Results for Different Task Sequences
@@
Recent work (Singh et al., 2022) has highlighted the consistent and substantial curriculum effects in class-incremental learning and continual visual question-answering tasks. Inspired by these findings, we conducted experiments to assess the impact of the curricula in the context of CSEGG. To delve into its potential influence, we trained baselines with three distinct task sequences for Learning Scenario 1. Our results demonstrated that curriculum learning indeed shapes CSEGG performance within class-incremental settings. Notably, in Fig. S13(d), a large difference in Avg.R@20 between Order 1 and Order 3 emerges for the Replay (100%) baseline. Similarly, Fig. S13(a) reveals a substantial Avg.R@20 disparity between Order 2 and Order 3 for the Naive baseline. This trend extends to F@20, as depicted in Fig. S13(e)(f)(g)(h). These insights collectively affirm the significance of the curricula within CSEGG.
@@
Figure S13: Results of F@K=20, Avg. R@K=20 over tasks on CSEGG models with the SGTR backbone in Learning Scenario 1 with different permutations of task sequences. (a),(e) denotes Avg.R@K and F@K for naive baseline. (b),(f) denotes Avg.R@K and F@K for EWC baseline. (c),(g) denotes Avg.R@K and F@K for Replay(10%) baseline. (d),(h) denotes Avg.R@K and F@K for Replay(100%) baseline. See Sec. 3.3 for introduction to continual learning baselines. See Sec. 3.4 for explanations about evaluation metrics. X-axis indicates the task numbers. The higher F, Avg.R@20, the better.
A.4.5 Minimal Forgetting in DETR
@@
To validate the impact of fine-tuning the DETR model in training Stage 2 of learning scenario S1 on relationship predicate predictions and to ensure minimal forgetting occurs in object detection (Sec 3.2), we compare the mean Average Precision (mAP) for object detection on the entire test set of Visual Genome between the pre-trained DETR checkpoint from the paper (Li et al., 2022b) and the DETR models after fine-tuning on each task of S1.
@@
Figure S14: mAP performance on the entire Visual Genome Test Set. We used a pre-trained DETR checkpoint (Li et al., 2022b) as well as a naive baseline of SGTR continually trained on each task from S1, evaluated them on the entire Visual Genome’s test set, and reported their mAP performances. We observed a minimal decrease in mAP across all tasks.
A.5 Comparision between CNN-SGG Backbone and the SGTR Model
[…]
Figure S15: Results of average recall and forgetting over tasks in Learning Scenario 1 (a and b) for CNN-SGG, when K-20. See Sec. 3.3 for introduction to continual learning baselines. See Sec. 3.4 for explanations about evaluation metrics. X-axis indicates the task numbers. The higher Avg.R@20 and F@20, the better.
Figure S16: Results of average recall and forgetting over tasks in Learning Scenario 2 (a and b) for CNN-SGG (K=20). See Sec. 3.3 for introduction to continual learning baselines. See Sec. 3.4 for explanations about evaluation metrics. X-axis indicates the task numbers. The higher Avg.R@20 and F@50, the better.
Fig. S17 provides the generalization results in detecting unknown objects and classifying known relationships among these objects in Learning Scenario 3 (S3) for CNN-SGG backbone. As elaborated in Sec. 4.3, an intriguing trend emerges: the CSEGG models demonstrate an increasing ability to identify unknown objects as the number of tasks rises. Furthermore, similar to our observations in Sec. 4.3, the CNN-SGG model exhibits proficiency in identifying known relationships between unfamiliar objects. However, this capability reaches a plateau as the task count escalates.
Figure S17: Generalization results in Learning Scenario 3 for CNN-SGG (K=20). See Sec. 3.4 for evaluation metrics. The higher the values, the better. Line colors indicate continual learning baselines. Line types denote the IoU thresholds for determining correctly predicted bounding box locations.
@@ A.5 Comparision between CNN-SGG Backbone and the SGTR Model
A.6 Visualization Examples for All Learning Scenarios for Continual SGTR-based Models
@@
In this section, we present visualization examples from each learning scenario to showcase the performance of the three continual SGTR-based models, namely Replay10%, EWC, and Naive, in three learning scenarios.
A.6.1 Learning Scenario 1 (S1)
@@
Figure S18: Visualization example for Learning Scenario 1 (S1). The leftmost column in the figure displays the ground truth bounding boxes and scene graphs for each task in Learning Scenario 1 (S1). The remaining columns, from left to right, represent the bounding boxes and scene graphs generated by each baseline model (Replay 10%, Naive, and EWC). In all the scene graphs, red boxes indicate objects, while green boxes represent relationships. The direction of the arrows between the red (object) and green (relationship) boxes indicates the subject and object ordering in the triplet. For example, in the scene graph predicted by the EWC model after Task 5, the triplet is "trunk behind women", as the arrow goes from "trunk" to "behind" to "women". The time arrow on the left side of the figure demonstrates that the model is exposed to new data over time, with new relationships incrementally added, as described in Sec. 3.1.
@@
Figure S19: Visualization example for Learning Scenario 2 (S2). The leftmost column shows the ground truth bounding boxes and scene graphs in each task of Learning Scenario 2 (S2). The remaining columns, from left to right, represent the bounding boxes and scene graphs generated by each baseline model (Replay 10%, Naive, and EWC). In all the scene graphs, red boxes indicate objects, while green boxes represent relationships. As explained in Fig. S18 caption, the direction of the arrows between the red (object) and green (relationship) boxes indicates the subject and object ordering in the triplet. The time arrow on the left side of the figure demonstrates that the model is exposed to new data over time, with new objects and relationships incrementally added, as described in Sec. 3.1.
@@
From Fig.S18 we observe that, in Task 1, the ground truth scene graph contains triplets of "on" relationship: "plate on table" and "hair on women". After training on task 1, all three models (Replay 10%, EWC, Naive) can accurately predict these triplets of "on" relationship.
@@
In Task 2, triplets of "has" relationship are introduced: "plate has food" and "women has hair". After training on task 2 data, the Replay 10% model successfully remembers the triplets of "on" relationship ("plate on table", "hair on women") from Task 1 and predicts "women has hair". The Naive model forgets the triplets of "on" relationship and only predicts "women has hair". The EWC model remembers the triplets of "on" relationships and predicts "women has hair". None of the models predict "plate has food".
@@
In Task 3, triplet of "at" relationship is introduced: "women at table". After training on task 3 data, the Replay 10% model remembers the previous triplets ("plate on table", "women has hair", "hair on women") and predicts "women at table". The Naive model forgets the previous triplets and only predicts "women at table". In contrast to its previous performance, the EWC model forgets the previous triplets and only predicts "women at table".
@@
In Task 4, triplets related to "of" relationship are introduced: "hand of women", "neck of women", "arm of women", and "head of women". After training on task 4 data, the Replay 10% model remembers the triplets related to "on" and "has" relationships ("plate on table", "women has hair", "hair on women") from previous tasks but forgets the "at" relationship triplet (“women at table”). It only predicts "head of women" from the triplets introduced in Task 4. The Naive and EWC models both forget the "at" relationship triplet from the previous task but predict "head of women" and "arm of women" from the triplets introduced in Task 4.
@@
In Task 5, triplet belonging to the "behind" relationship is introduced: "trunk behind women". After training on task 5 data, the Replay 10% model forgets the triplets related to "on" relationship ("plate on table", "hair on women") and only remembers the triplets related to "has" and "of" relationships ("women has hair", "head of women") learned from the previous task. It is not able to predict "trunk behind women". The Naive model, similar to its performance after previous tasks, fails to remember any triplets previously learned and only predicts "trunk behind women". The EWC model also fails to remember any triplets from the previous task and only predicts "trunk behind women".
A.6.2 Learning Scenario 2 (S2)

### v3

@@ 0.A.3 Evaluation Metrics
To assess whether the knowledge at previous tasks facilitates learning the new task and whether the knowledge at new tasks enhances the performances at older tasks, we introduce Forward Transfer (FWT@K) [34] and Backward Transfer (BWT@K)[35]. BWT@K is defined as $BWT@K=\frac{1}{T-1}\sum_{i=1}^{T-1}R@K_{T,i}\>-\>R@K_{i,i}$, where $T$ denotes the total number of tasks in a learning scenario and $R@K_{i,j}$ denotes the continual learning model trained after task $i$ and tested in task $j$. FWT is defined as $FWT@K=\frac{1}{T-1}\sum_{i=2}^{T}R@K_{i,i}\>-\>\overline{b@K}_{i,i}$, where $\overline{b@K}_{i,i}$ is the R@K for an independent model with random initialization trained in task $i$ and tested in task $i$.
In learning scenario S3, we evaluate CSEGG models on their abilities to generalize to detect unknown objects and classify known relationships on these objects, in the standalone generalization test set over all tasks. To benchmark these, we introduce two evaluation metrics: the recall of the predicted bounding boxes on unknown objects (Gen Rbbox@K) and the recall of the predicted graph $G_{i}$ (Gen R@K). As the CSEGG models have never been taught to classify unknown objects, we discard the class labels of the bounding boxes and only evaluate the predicted box locations with Gen Rbbox@K. To evaluate whether the predicted box location is correct, we apply a hard threshold of Intersection over Union (IoU) between the predicted bounding box locations and the ground truth. Any predicted bounding boxes with their IoU values above the hard threshold are deemed to be correct. We vary IoU thresholds from 0.3, 0.5, and 0.7. To assess whether the CSEGG model generalizes to detect known relationships over unknown objects, we evaluate the recall Gen R@K of the predicted relationships $r_{k}$ only on correctly predicted bounding boxes.
0.A.4 More Quantitative Result Analysis on Continual SGTR Methods
0.A.4.1 Results for Different Task Sequences
@@
Recent work [54] has highlighted the consistent and substantial curriculum effects in class-incremental learning and continual visual question-answering tasks. Inspired by these findings, we conducted experiments to assess the impact of the curricula in the context of CSEGG. To delve into its potential influence, we trained baselines with three distinct task sequences for Learning Scenario 1. Our results demonstrated that curriculum learning indeed shapes CSEGG performance within class-incremental settings. Notably, in Fig. S3(d), a large difference in Avg.R@20 between Order 1 and Order 3 emerges for the Replay (100%) baseline. Similarly, Fig. S3(a) reveals a substantial Avg.R@20 disparity between Order 2 and Order 3 for the Naive baseline. This trend extends to F@20, as depicted in Fig. S3(e)(f)(g)(h). These insights collectively affirm the significance of the curricula within CSEGG.
@@
Figure S3: Results of F@K=20, Avg. R@K=20 over tasks on CSEGG models with the SGTR backbone in Learning Scenario 1 with different permutations of task sequences. (a),(e) denotes Avg.R@K and F@K for naive baseline. (b),(f) denotes Avg.R@K and F@K for EWC baseline. (c),(g) denotes Avg.R@K and F@K for Replay@10% baseline. (d),(h) denotes Avg.R@K and F@K for Replay@100% baseline. See Sec. 3.2 for introduction to continual learning baselines. See Sec. 3.3 for explanations about evaluation metrics. X-axis indicates the task numbers. The higher F, Avg.R@20, the better.
0.A.4.2 Minimal Forgetting in DETR
@@
To validate the impact of fine-tuning the DETR model in training Stage 2 of learning scenario S1 on relationship predicate predictions and to ensure minimal forgetting occurs in object detection (Sec 0.A.2), we compare the mean Average Precision (mAP) for object detection on the entire test set of Visual Genome between the pre-trained DETR checkpoint from the paper [28] and the DETR models after fine-tuning on each task of S1.
@@
Figure S4: mAP performance on the entire Visual Genome Test Set. We used a pre-trained DETR checkpoint [28] as well as a naive baseline of SGTR continually trained on each task from S1, evaluated them on the entire Visual Genome’s test set, and reported their mAP performances. We observed a minimal decrease in mAP across all tasks.
@@
0.A.5 Visualization Examples for All Learning Scenarios for Continual SGTR-based Models
@@
In this section, we present visualization examples from each learning scenario to showcase the performance of the three continual SGTR-based models, namely Replay@10%, EWC, and Naive, in three learning scenarios.
0.A.5.1 Learning Scenario 1 (S1)
@@
Figure S5: Visualization example for Learning Scenario 1 (S1). The leftmost column in the figure displays the ground truth bounding boxes and scene graphs for each task in Learning Scenario 1 (S1). The remaining columns, from left to right, represent the bounding boxes and scene graphs generated by each baseline model (Replay@10%, Naive, and EWC). In all the scene graphs, red boxes indicate objects, while green boxes represent relationships. The direction of the arrows between the red (object) and green (relationship) boxes indicates the subject and object ordering in the triplet. For example, in the scene graph predicted by the EWC model after Task 5, the triplet is "trunk behind women", as the arrow goes from "trunk" to "behind" to "women". The time arrow on the left side of the figure demonstrates that the model is exposed to new data over time, with new relationships incrementally added, as described in Sec. 3.1.
@@
Figure S6: Visualization example for Learning Scenario 2 (S2). The leftmost column shows the ground truth bounding boxes and scene graphs in each task of Learning Scenario 2 (S2). The remaining columns, from left to right, represent the bounding boxes and scene graphs generated by each baseline model (Replay@10%, Naive, and EWC). In all the scene graphs, red boxes indicate objects, while green boxes represent relationships. As explained in Fig. S5 caption, the direction of the arrows between the red (object) and green (relationship) boxes indicates the subject and object ordering in the triplet. The time arrow on the left side of the figure demonstrates that the model is exposed to new data over time, with new objects and relationships incrementally added, as described in Sec. 3.1.
@@
From Fig.S5 we observe that, in Task 1, the ground truth scene graph contains triplets of "on" relationship: "plate on table" and "hair on women". After training on task 1, all three models (Replay@10%, EWC, Naive) can accurately predict these triplets of "on" relationship.
@@
In Task 2, triplets of "has" relationship are introduced: "plate has food" and "women has hair". After training on task 2 data, the Replay@10% model successfully remembers the triplets of "on" relationship ("plate on table", "hair on women") from Task 1 and predicts "women has hair". The Naive model forgets the triplets of "on" relationship and only predicts "women has hair". The EWC model remembers the triplets of "on" relationships and predicts "women has hair". None of the models predict "plate has food".
@@
In Task 3, triplet of "at" relationship is introduced: "women at table". After training on task 3 data, the Replay@10% model remembers the previous triplets ("plate on table", "women has hair", "hair on women") and predicts "women at table". The Naive model forgets the previous triplets and only predicts "women at table". In contrast to its previous performance, the EWC model forgets the previous triplets and only predicts "women at table".
@@
In Task 4, triplets related to "of" relationship are introduced: "hand of women", "neck of women", "arm of women", and "head of women". After training on task 4 data, the Replay@10% model remembers the triplets related to "on" and "has" relationships ("plate on table", "women has hair", "hair on women") from previous tasks but forgets the "at" relationship triplet (“women at table”). It only predicts "head of women" from the triplets introduced in Task 4. The Naive and EWC models both forget the "at" relationship triplet from the previous task but predict "head of women" and "arm of women" from the triplets introduced in Task 4.
@@
In Task 5, triplet belonging to the "behind" relationship is introduced: "trunk behind women". After training on task 5 data, the Replay@10% model forgets the triplets related to "on" relationship ("plate on table", "hair on women") and only remembers the triplets related to "has" and "of" relationships ("women has hair", "head of women") learned from the previous task. It is not able to predict "trunk behind women". The Naive model, similar to its performance after previous tasks, fails to remember any triplets previously learned and only predicts "trunk behind women". The EWC model also fails to remember any triplets from the previous task and only predicts "trunk behind women".
0.A.5.2 Learning Scenario 2 (S2)